OpenStudy (anonymous):

Can anyone provide me with some guidance on this problem? What is P(ABC) if P(a)=P(B)=P( C)=.5, P(AUB)=.55, P(AUC)=.7, P(BC)=.3, and P(ABC)=2P(ABC^(complement))?

OpenStudy (anonymous):

i have tried calculating some of the other things as well but I dont think they got me any closer to an answer. I also know the answer should be .3 but i dont understand how to get to that

OpenStudy (amistre64):

hmm, auc = a + c - ac aub = a + b - ab bc = b (c|b) hmm

OpenStudy (anonymous):

ya i found P(ab) to be .45 and P(AC) to be .3

OpenStudy (amistre64):

auc = a + c - ac = .7 aub = a + b - ab = .55 and a=b=c=.5 auc = .5 + .5 - ac = .7 aub = .5 + .5 - ab = .55

OpenStudy (anonymous):

using the first two you said

OpenStudy (amistre64):

ac = .3 ab = .45

OpenStudy (amistre64):

bc = .5 (c|b) = .3 c|b = 3/5

OpenStudy (anonymous):

i also know my professor said that this may be useful: P(AUBUC) = P(A)+P(B)+P(C)-P(AB)-P(AC)-P(BC)+P(ABC)

OpenStudy (amistre64):